at what x-coordinate does f(x)=|-2x+3|-2 have its minimum value?
a. -2
b. -2/3
c. 3/2
d. 3

Answer by stanbon(75887)

at what x-coordinate does f(x)=|-2x+3|-2 have its minimum value?
---------------------
y is minimum when |-2x+3| = 0
---
Solve -2x+3 = 0
-2x = -3
x = 3/2
----------------
y is at a minimum when x = 3/2
----------------
Cheers,
Stan H.